HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THEORETICAL DFT INSIGHTS INTO THE NATURE OF CHALCOGEN

BONDING BETWEEN SO_{2}, SO_{3} AND DIAZINE: NONCOVALENT INTERACTIONS INVOLVING π HOLES

\underline{Y. ARFAOUI}, , Laboratory of Caracterisations, Applications and Modeling Materials (LR18ES08), Department of Chemistry, Faculty of Science, University of

Tunis El Manar, Tunis.

The irritant gases SO_{2} and SO_{3} can be the cause of certain health problems, especially in sensitive people such as children. Short exposures can cause bronchial spasms, coughing spells, impaired respiratory function and eye irritation. By attaching to fine particles, a small part of this pollutant can even penetrate to the lower respiratory tract. Too high concentrations of SO_{2} and SO_{3} can cause acidification of the environment. Once emitted into the air and in the presence of water, SO_{2} and SO_{3} are transformed into sulfuric acid H_{2}SO_{4} which contributes to the phenomenon of acid rain. The consequences of such rains are acid deposits on building materials which can cause damage to the architectural and cultural heritage, the acidification of surface waters leading to the death of fish and soil deterioration which can have harmful effects on the vegetation and plants. DFT calculation of the capture of SO_{2} and SO_{3} viaπholes is performed with diazine molecules. The charge transfer interactions is highlighted by the wiberg indices calculation and the NBO analysis. The IRC calculation is performed to connect the reactants SO_{2} and SO_{3} with the H_{2}SO_{4} product via the transition states.

p-number: p003  Submitted on Thu, 11 May 2023 11:05:03 +0200